# A Tale of Single-channel Electroencephalogram: Devices, Datasets, Signal Processing, Applications, and Future Directions

### v1

@@ IV DATASETS of SINGLE-CHANNEL EEG
The K-EmoCon dataset [40] is a multimodal sensor dataset designed for continuous emotion recognition in natural conversations. It addresses the limitations of existing emotion datasets in capturing individual emotions in natural environments. K-EmoCon includes audio and video recordings, EEG and peripheral physiological signals collected from 32 participants during 16 debates, each approximately 10 minutes long. The EEG signals were collected using the NeuroSky MindWave headset, which uses two dry sensor electrodes: one on the forehead (FP1 channel of the frontal lobe - 10/20 system) and the other on the left earlobe (reference). The data collection process consists of four stages: entry, baseline measurement, debate and emotion annotation. While watching the debate video, participants annotated their emotions every 5 seconds, including self-annotation, opponent annotation and external observer annotation. The annotations cover arousal-valence and 18 emotion categories. The data has been thoroughly checked and preprocessed to ensure signal integrity and device reliability.
@@
Acampora et al. proposed a single-channel EEG dataset from a portable steady-state visual evoked potential (SSVEP) based brain-computer interfaces (BCI) involving 11 volunteers [38]. The data were acquired by OlimexEEG-SMT, a dual-channel 10-bit analog-to-digital converter with a sampling frequency of 256 Hz. The experiment utilized the repetitive visual stimuli with four flickering frequencies: 8.57 Hz, 10 Hz, 12 Hz, and 15 Hz, and each lasting for 16 seconds. Only one channel was used, with two active electrodes (CH- and CH+) placed at Fpz (frontal region) and Oz (occipital region), and the reference electrode on the earlobe. This dataset serves as a crucial baseline for single-channel EEG, particularly in the emerging field of human-robot interaction.
The scarcity of single-channel EEG datasets poses a significant challenge in current neuroscience research. Expanding unipolar single-channel EEG datasets to obtain more reliable data is essential. Compared to multi-channel EEG datasets, unipolar single-channel EEG datasets have obvious advantages in equipment cost, data processing complexity and practical application [42]. However, expanding these datasets requires addressing challenges in data collection, annotation and data quality control.
TABLE II: Single-channel EEG Datasets
  Datasets EEG-Devices Collected Data Sampling Rate Signal Range[min,max]
K-EmoCon [40] NeuroSky MindWave Headset Brainwave (Fp1 Channel EEG) 125HZ n/a
Attention & Meditation 1HZ [0,100]
Emo_Food [41] NeuroSky MindWave Mobile II Brainwave (Fp1 Channel EEG) 512HZ n/a
SSVEP_BCI [38] Olimex EEG-SMT Brainwave (Fpz or Oz Channel EEG) 256HZ n/a
@@ V SIGNAL PROCESSING of SINGLE-CHANNEL EEG
In this section, we present a detailed processing framework of single-channel EEG signals, encompassing acquisition, preprocessing, feature extraction, feature selection and machine learning or deep learning-based downstream tasks, as depicted in Fig4.
@@ V-A1 Basic Preprocessing & Filtering
In the preprocessing stage of single-channel EEG signals, removing baseline drift and power supply interference is a key pre-processing step. Baseline drift in EEG signals refers to a type of low-frequency signal variation caused by poor electrode contact or other slowly changing factors. High-pass filtering or empirical mode decomposition (EMD) methods are usually used to remove low-frequency drifts, including baseline drifts [43]. To remove power supply interference, notch filters or adaptive filtering are often applied to specifically eliminate power frequency interference of 50Hz or 60Hz [44]. Based on the designed filters, single-channel EEG signals are usually divided into five different frequency bands: $\delta$ (0.5-4 Hz), $\theta$ (4-7 Hz), $\alpha$ (8-13 Hz), $\beta$ (14-30 Hz), and $\gamma$ (>30, but usually <100 Hz).
@@ V-A2 Artifacts Cancellation
Single-channel EEG artifact cancellation faces critical challenges, including the broad spectral impact of artifacts, the lack of standardized preprocessing steps that effectively filter noise while preserving signal integrity and high variability among subjects and trials in phase-locked EEG signals. These artifacts mainly include electrooculogram (EOG) artifacts, electromyography (EMG) artifacts, and motion artifacts [45]. EOG artifacts caused by blinking and eye movement appear as low-frequency, high-amplitude signals [46], while EMG artifacts result from scalp and facial muscle activity and are characterized by a wide frequency range and high amplitude [47]. Motion artifacts arise from poor electrode contact due to head or body movement, producing low-frequency, large-amplitude noise [48]. These artifacts can significantly reduce the accuracy of single-channel EEG data, increase the risk of misdiagnosis and complicate data processing. For instance, eye movement artifacts may be mistaken for epileptic seizures, while EMG artifacts may obscure genuine EEG abnormalities, leading to missed or incorrect diagnosis. These artifacts also limits the extensive application of single-channel EEG in portable and mobile devices, particularly in scenarios requiring high-precision.
@@
Since single-channel EEG systems lack multi-channel information commonly used in multi-channel EEG artifact removal methods such as independent component analysis (ICA), canonical correlation analysis (CCA) and independent vector analysis (IVA) [45], they cannot be directly applied. Some studies have applied decomposition methods such as ensemble empirical mode decomposition (EEMD) [49] and singular spectrum decomposition (SSA) [50] to convert single-channel EEG signals into a multi-dimensional format. Hybrid methods, such as EEMD-ICA [51] and SSA-ICA [52, 53], have also been employed to process single-channel EEG. EEMD-ICA first decomposes the signal into spectral independent modes using EEMD, then applies ICA to extract statistically independent sources [51]. In [52, 53], SSA decomposes a single-channel EEG signal into multiple sub-signals to form a multi-channel dataset, and ICA extracts statistically independent source signals from these sub-signals to identify and remove artifact-containing components. To avoid the edge effect in EEMD-ICA, Guo et al. proposed combining local mean decomposition (LMD) and ICA (LMD-ICA) [54], although parameter adjustment in LMD is challenging. [55] proposed a combination of discrete wavelet transform (WT) and w-ICA for denoising EEG signals [56]. In many biomedical applications based on single-channel EEG, artifacts often appear in narrow frequency bands. Thus, hybrid methods including ICA are more suitable for removing stereotyped artifacts, although ICA-based hybrid methods do not perform well due to the significant variations in multiple types of artifacts [57] and will shorten the detection time and reduce the battery life because of the high computational cost, which results in increased power consumption.
In contrast, the low computational complexity of CCA extends the working time of single-channel EEG devices. A technique combining EEMD and CCA was proposed in [58, 59] to eliminate motion artifacts in single-channel EEG signals. This technique first uses EEMD to decompose the single-channel signal into a set of oscillatory components (intrinsic mode functions, IMFs), converting it into a multivariate signal. CCA then extracts the source signal from the multivariate data and its delayed version. The choice of method depends on the artifact type: ICA-based methods are typically used for EOG artifacts, while CCA and IVA are better suited for EMG artifacts. Currently, no single method can completely remove all types of artifacts.
@@
Compared with CCA, WT is a lower computational complexity method that can be used. WT represents the time-frequency pattern of a signal by decomposing it into multiple groups of coefficients. Compared to blind source separation, WT has a lower computational cost, making it more suitable for online applications of single-channel EEG. In [60], a method combining discrete wavelet transform (DWT) and adaptive noise cancellation (ANC) was developed to remove eye movement artifacts (OA) from EEG. Khatun et al. systematically evaluated the effectiveness of unsupervised WT decomposition techniques, specifically discrete DWT and stationary wavelet transform (SWT), using four wavelet basis functions (haar, coif3, sym3, and bior4.4) and two thresholding methods (Universal Threshold and Statistical Threshold) for ocular artifact removal from single-channel EEG data [61].
Moreover, Tengtrairat et al. proposed an online adaptive single-channel source separation method using a pseudo-stereo noisy mixing model and an adaptive spectrum amplitude estimator to enhance and separate mixtures of sources from noisy environments [62]. Noorbasha and Sudha integrated overlap segmentation, adaptive singular spectrum analysis (ASSA), and ANC to remove EOG artifacts from single-channel EEG [63]. Shahbakhti et al. proposed the VME-DWT algorithm to detect and filter eye blinks from short segments of single-channel EEG [64]. Noorbasha and Sudha combined SSA, ICA, and SWT to remove EOG artifacts and separate cerebral activity components from single-channel EEG signals [65].
@@
In addition, deep learning (DL) has been proposed to effectively capture artifacts in EEG signals. Lu et al. introduced Dual-Stream Attention-TCN for removing EMG artifacts from single-channel EEG signals [66]. This approach employs two parallel streams processing EEG signals in different frequency bands: the first stream focuses on extracting low-frequency EEG features to mitigate EMG interference, while the second stream integrates high-level features from the first stream with its own low-level features for comprehensive EEG reconstruction across the frequency spectrum. Cui et al. proposed EEGIFNet, a dual-branch interactive fusion neural network, for artifact removal in single-channel EEG signals [67]. EEGIFNet predicts clean EEG signals and residual artifacts through two branches, facilitating information exchange and fusion via interactive and fusion modules to effectively eliminate various types of artifacts. Zhang et al. presented a two-stage intelligent multi-type artifact removal model based on GRU autoencoder. This model aims to remove mixed multi-type artifacts from single-channel EEG by first identifying and then removing artifacts [68]. It employs attention mechanisms and adaptive feature extraction in the encoding feature domain to distinguish contaminated and clean EEG signals, achieving efficient artifact removal.
@@
Figure 4: Flowchart of single-channel EEG signal processing illustrating key steps from signal acquisition to downstream tasks, encompassing preprocessing, feature extraction, feature selection, as well as traditional machine learning (ML) and deep learning (DL) methods.
@@ V-B Feature Extraction
Time domain feature extraction focuses on deriving information directly from the EEG signals. These features encompass statistical properties such as mean, variance, skewness, and kurtosis, alongside characteristics like amplitude, duration, and waveform dynamics of the signal. This approach is advantageous as it captures transient behaviors and dynamic traits within EEG signals. In addition to fundamental statistical features, more intricate aspects can be revealed through pattern recognition techniques applied to time series data, such as tracking changes in frequency components over time using methods like short-time Fourier transform (STFT) [69, 70].
@@ V-C Feature Selection
Feature selection is a crucial step in EEG signal processing, aimed at identifying and selecting the most informative and relevant features from high-dimensional datasets. Given the high noise levels and large volumes of EEG data, effective feature selection is essential for accurately interpreting brain activity. Techniques range from traditional statistical tests like the t-test and analysis of variance to more complex methods such as wrapper and embedded techniques [73, 74]. Innovations in deep learning, including autoencoders and tree-based models, further improve feature selection by automatically identifying and quantifying the importance of EEG features [75]. For a detailed overview, refer to previous reviews [76, 77].
@@ V-D Machine Learning of Single-channel EEG
In the processing of single-channel EEG signals, machine learning (ML) has significantly advanced automation and precision in targeted downstream tasks [78]. Traditional ML methods typically involve manual extraction and selection of features such as time domain statistical features, frequency domain features, and time-frequency domain features. These features are subsequently utilized to train various ML models like support vector machines (SVM), random forests (RF), and k-nearest neighbors (KNN) for specific tasks. While these models have demonstrated effective performance in EEG signal classification, their accuracy heavily relies on the quality and relevance of the chosen features [79, 80, 81].
Deep learning (DL) techniques have shown significant efficacy in EEG signal processing by leveraging multi-layer nonlinear architectures to learn intricate data features [82, 83]. Unlike traditional ML methods, DL can automatically derive complex feature representations from preprocessed data without manual feature engineering. For instance, convolutional neural networks (CNNs) excel in processing one-dimensional sequence data and have been successfully applied to automatic EEG feature extraction [84]. Recurrent neural networks (RNNs) and long short-term memory networks (LSTMs) effectively handle the time series characteristics of EEG signals and capture temporal dependencies [82]. These capabilities have enabled DL models to achieve significant progress in tasks such as EEG signal classification, anomaly detection, and pattern recognition. Although DL excels in feature learning, it typically requires extensive training data and has poor interpretability. Conversely, traditional ML methods can operate with smaller datasets and offer more transparent decision-making processes. Therefore, selecting an appropriate ML strategy in practical applications should consider data availability, task complexity, and the need for model interpretability.
@@ VI-A Sleep Staging
The Fpz-Cz channel has been extensively utilized in sleep studies due to its efficiency in detecting wakefulness and REM stages, as well as its ease of use [93, 94, 95]. Wei et al. employed Fpz-Cz channel EEG data from the Sleep-EDFx database, combining contextual scale maps and Co-ScaleNet to achieve high accuracy and recall in automatic sleep staging [93]. Their method proved robust across multiple datasets, including those from healthy individuals, patients with depression and other public datasets, particularly in the N1 stage.
Zhang et al. utilized Fpz-Cz channel EEG data from the Sleep-EDFx dataset and proposed a semi-supervised learning model, SHNN, for automatic sleep staging. This model demonstrated high classification accuracy and consistency, especially in light and deep sleep stages [21]. Studies by [96, 22] used single-channel EEG data from the Fpz-Cz channel for automated sleep staging, exploring sleep staging directly from EEG signals using temporal context information. Eldele et al. proposed an attention-based DL architecture, AttnSleep, to classify sleep stages using single-channel EEG signals, whose architecture relies on a multi-resolution CNN and adaptive feature recalibration [84]. In the context of sleep stage scoring on the Fpz-Cz channel, [82] developed an efficient single-channel EEG analysis model based on CNN and LSTM. Seifpour et al. proposed an automatic sleep staging system based on local extreme value statistical behavior, utilizing statistical methods to extract features from the Fpz-Cz channel. Qu et al. developed a method based on the residual attention model, achieving accurate EEG sleep staging using DL technology on the Fpz-Cz channel [97, 98].
The C4-A1 and Cz-A1 channels have demonstrated effective performance for sleep staging tasks on certain datasets. In the SHHS dataset, the C4-A1 channel shows a clear advantage by capturing more information when processing complex EEG signals, thereby improving classification accuracy [96, 99]. The Cz-A1 channel, known for providing stable signals, helps reduce misclassification due to eye movements and its application across multiple datasets has yielded consistent and high-precision classification results.
[97] used the same local extreme value statistical method on the C4-A1 channel for automatic sleep staging and verified the validity of the C4-A1 channel signal. [84] also conducted further experiments on the Cz-A1 channel signal and obtained excellent sleep stage classification results, highlighting the advantages of the Cz-A1 channel in capturing stable signals and reducing eye movement interference. Furthermore, Supratak and Guo proposed the TinySleepNet model, which employs DL technology for efficient single-channel EEG analysis. Their study utilized the Cz-A1 channel to verify its efficiency and versatility across multiple datasets. The TinySleepNet model achieves performance comparable to or better than existing methods while reducing the number of parameters and computing resources required [82].

### v2

@@ IV DATASETS of SINGLE-CHANNEL EEG
The K-EmoCon dataset [61] is a multimodal sensor dataset designed for continuous emotion recognition in natural conversations. It addresses the limitations of existing emotion datasets in capturing individual emotions in natural environments. K-EmoCon includes audio and video recordings, EEG and peripheral physiological signals collected from 32 participants during 16 debates, each approximately 10 minutes long. The EEG signals were collected using the NeuroSky MindWave headset, which uses two dry sensor electrodes: one on the forehead (FP1 - 10/20 system) and the other on the left earlobe (reference). The data collection process consists of four stages: entry, baseline measurement, debate and emotion annotation. While watching the debate video, participants annotated their emotions every 5 seconds, including self-annotation, opponent annotation and external observer annotation. The annotations cover arousal-valence and 18 emotion categories. The data has been thoroughly checked and preprocessed to ensure signal integrity and device reliability.
@@
Acampora et al. proposed a single-channel EEG dataset from a portable steady-state visual evoked potential (SSVEP) based BCI involving 11 volunteers [49]. The recorded EEG data from a single volunteer contains the response to an intermittent source of light, which is emitted at four different frequencies, namely 8.57 Hz (F1), 10 Hz (F2), 12 Hz (F3) and 15 Hz (F4). At the end of the experiment, a total of 44 recordings of 16 s each one sampled at 256 Hz, for a total of 180224 samples was obtained. This dataset serves as a crucial baseline for single-channel EEG, particularly in the emerging field of human-robot interaction.
The scarcity of single-channel EEG datasets poses a significant challenge in current neuroscience research, and expanding these datasets is essential given their advantages over multi-channel EEG datasets in equipment cost, data processing complexity and practical application [63]. However, expanding these datasets requires addressing challenges in data collection, annotation and data quality control.
TABLE II: Single-channel EEG Datasets with Additional Details
  Dataset EEG Device Collected Data Sampling Rate Subjects Signal Duration
K-EmoCon [61] NeuroSky MindWave Headset Brainwave (Fp1 Channel EEG) 125 Hz 32 Approx. 10 min per session
Attention & Meditation 1 Hz
Emo_Food [62] NeuroSky MindWave Mobile II Brainwave (Fp1 Channel EEG) 512 Hz 20 3 trials per subject (duration not specified)
SSVEP_BCI [49] Olimex EEG-SMT Brainwave (Fpz or Oz Channel EEG) 256 Hz 11 Approx. 64 s per subject
@@ V SIGNAL PROCESSING of SINGLE-CHANNEL EEG
In this section, we present a detailed processing framework of single-channel EEG signals, encompassing acquisition, preprocessing, feature extraction, feature selection and ML or DL-based downstream tasks, as depicted in Fig. 4.
@@ V-A1 Basic Preprocessing & Filtering
In the preprocessing stage of single-channel EEG signals, removing baseline drift and power supply interference is a key pre-processing step. Baseline drift in EEG signals refers to a type of low-frequency signal variation caused by poor electrode contact or other slowly changing factors. High-pass filtering or empirical mode decomposition (EMD) methods are usually used to remove low-frequency drifts, including baseline drifts [64]. To remove power supply interference, notch filters or adaptive filtering are often applied to specifically eliminate power frequency interference of 50 Hz or 60 Hz [65]. Based on the designed filters, single-channel EEG signals are usually divided into five different frequency bands: $\delta$ (0.5-4 Hz), $\theta$ (4-7 Hz), $\alpha$ (8-13 Hz), $\beta$ (14-30 Hz), and $\gamma$ (>30, but usually <100 Hz) [66].
@@ V-A2 Artifacts Cancellation
Single-channel EEG artifact cancellation faces critical challenges, including the broad spectral impact of artifacts, the lack of standardized preprocessing steps that effectively filter noise while preserving signal integrity and high variability among subjects and trials in phase-locked EEG signals. These artifacts mainly include electrooculogram (EOG) artifacts, electromyography (EMG) artifacts, and motion artifacts [67]. EOG artifacts from blinking and eye movement appear as low-frequency, high-amplitude signals [68], EMG artifacts from scalp and facial muscle activity show wide frequency range and high amplitude [69], while motion artifacts due to poor electrode contact produce low-frequency, large-amplitude noise [70]. These artifacts can significantly reduce the accuracy of single-channel EEG data, increase the risk of misdiagnosis and complicate data processing. For instance, eye movement artifacts may be mistaken for epileptic seizures, while EMG artifacts may obscure genuine EEG abnormalities, leading to missed or incorrect diagnosis. These artifacts also limits the extensive application of single-channel EEG in portable and mobile devices, particularly in scenarios requiring high-precision.
@@
Since single-channel EEG systems lack multi-channel information commonly used in multi-channel EEG artifact removal methods such as independent component analysis (ICA), canonical correlation analysis (CCA) and independent vector analysis (IVA) [67], they cannot be directly applied. These methods often fall under the category of blind source separation (BSS) [71], a technique used to separate a set of mixed signals into their original components without prior knowledge of the source signals or the mixing process. Some studies have applied decomposition methods such as ensemble empirical mode decomposition (EEMD) [72] and singular spectrum analysis (SSA) [73] to convert single-channel EEG signals into a multi-dimensional format. Hybrid methods have also been employed, where EEMD-ICA [74] decomposes signals into spectral independent modes, while SSA-ICA [75, 76] decomposes signals into multiple sub-signals, both followed by ICA to extract statistically independent sources for artifact removal. To avoid the edge effect in EEMD-ICA, Guo et al. proposed combining local mean decomposition (LMD) and ICA (LMD-ICA) [77], although parameter adjustment in LMD is challenging. Inuso et al. proposed a method called wavelet-Independent Component Analysis (w-ICA) for EEG signal denoising that combines discrete wavelet transform (DWT) and ICA [78]. In many biomedical applications based on single-channel EEG, artifacts often appear in narrow frequency bands [79]. Thus, hybrid methods including ICA are more suitable for removing stereotyped artifacts, although ICA-based hybrid methods do not perform well due to the significant variations in multiple types of artifacts [80] and will shorten the detection time and reduce the battery life because of the high computational cost, which results in increased power consumption. In contrast, the low computational complexity of CCA extends the working time of single-channel EEG devices. A technique combining EEMD and CCA was proposed in [81, 82] to eliminate motion artifacts in single-channel EEG signals. This technique first uses EEMD to decompose the single-channel signal into a set of oscillatory components (intrinsic mode functions, IMFs), converting it into a multivariate signal. CCA then extracts the source signal from the multivariate data and its delayed version, with ICA-based methods typically used for EOG artifacts and CCA and IVA better suited for EMG artifacts, though no single method can completely remove all types of artifacts.
@@
Compared with CCA, wavelet transform (WT) is a lower computational complexity method that can be used. WT represents the time-frequency pattern of a signal by decomposing it into multiple groups of coefficients. Compared to BSS, WT has a lower computational cost, making it more suitable for online applications of single-channel EEG. A method combining DWT and adaptive noise cancellation (ANC) was developed to remove eye movement artifacts from EEG [83]. Khatun et al. systematically evaluated the effectiveness of unsupervised WT decomposition techniques, specifically DWT and stationary wavelet transform (SWT), using four wavelet basis functions (haar, coif3, sym3, and bior4.4) and two thresholding methods (Universal Threshold and Statistical Threshold) for ocular artifact removal from single-channel EEG data [84]. Tengtrairat et al. proposed an online adaptive source separation method using pseudo-stereo noisy mixing and adaptive spectrum amplitude estimation [85]. Subsequent studies integrated various techniques: overlap segmentation with adaptive singular spectrum analysis (ASSA) and ANC [86], Variational Mode Extraction-DWT (VME-DWT) algorithm for eye blink detection [87], and combination of SSA, ICA, and SWT for EOG artifact removal [76].
@@
In addition, DL has been proposed to effectively capture artifacts in EEG signals. Lu et al. introduced Dual-Stream Attention-TCN for removing EMG artifacts from single-channel EEG signals [88]. This approach employs two parallel streams processing EEG signals in different frequency bands: the first stream focuses on extracting low-frequency EEG features to mitigate EMG interference, while the second stream integrates high-level features from the first stream with its own low-level features for comprehensive EEG reconstruction across the frequency spectrum. Cui et al. proposed EEGIFNet, a dual-branch interactive fusion neural network, for artifact removal in single-channel EEG signals [89]. EEGIFNet predicts clean EEG signals and residual artifacts through two branches, facilitating information exchange and fusion via interactive and fusion modules to effectively eliminate various types of artifacts. Zhang et al. presented a two-stage intelligent multi-type artifact removal model based on GRU autoencoder. This model aims to remove mixed multi-type artifacts from single-channel EEG by first identifying and then removing artifacts [90]. It employs attention mechanisms and adaptive feature extraction in the encoding feature domain to distinguish contaminated and clean EEG signals, achieving efficient artifact removal.
While traditional signal processing approaches like BSS and wavelet-based methods have established foundations for artifact removal, the field has evolved toward hybrid architectures and DL solutions that offer enhanced adaptability and robustness. Nevertheless, the optimal balance between computational efficiency and removal accuracy remains a crucial consideration for practical applications, especially in resource-constrained portable devices.
@@
Figure 4: Flowchart of single-channel EEG signal processing illustrating key steps from signal acquisition to downstream tasks, encompassing preprocessing, feature extraction, feature selection, as well as traditional machine learning and deep learning methods.
@@ V-B Feature Extraction
Time domain feature extraction focuses on deriving information directly from the EEG signals [91]. These features include statistical properties (mean, variance, skewness, kurtosis) and signal characteristics (amplitude, duration, waveform dynamics), effectively capturing transient behaviors and dynamic traits within EEG signals. In addition to fundamental statistical features, more intricate aspects can be revealed through pattern recognition techniques applied to time series data, such as tracking changes in frequency components over time using methods like short-time Fourier transform (STFT) [92, 93].
@@ V-C Feature Selection
Feature selection is a crucial step in EEG signal processing, aimed at identifying and selecting the most informative and relevant features from high-dimensional datasets. Given the high noise levels and large volumes of EEG data, effective feature selection is essential for accurately interpreting brain activity. Techniques range from traditional statistical tests like the t-test and analysis of variance to more complex methods such as wrapper and embedded techniques [96, 97]. Innovations in DL, including autoencoders and tree-based models, further improve feature selection by automatically identifying and quantifying the importance of EEG features [98]. For a detailed overview, refer to previous reviews [99, 42].
@@ V-D Machine Learning of Single-channel EEG
In the processing of single-channel EEG signals, ML has significantly advanced automation and precision in targeted downstream tasks [100]. Traditional ML methods typically involve manual extraction and selection of features such as time domain statistical features, frequency domain features, and time-frequency domain features. These features are subsequently utilized to train various ML models like support vector machines (SVM), random forests (RF), and k-nearest neighbors (KNN) for specific tasks. While these models have demonstrated effective performance in EEG signal classification, their accuracy heavily relies on the quality and relevance of the chosen features [101, 102, 103].
The DL techniques have shown significant efficacy in EEG signal processing through multi-layer nonlinear architectures that automatically learn intricate data features and derive complex feature representations without manual engineering [104, 105], unlike traditional ML methods. For instance, convolutional neural networks (CNNs) excel in processing one-dimensional sequence data and have been successfully applied to automatic EEG feature extraction [106]. Recurrent neural networks (RNNs) and long short-term memory networks (LSTMs) effectively handle the time series characteristics of EEG signals and capture temporal dependencies [104]. These capabilities have enabled DL models to achieve significant progress in tasks such as EEG signal classification, anomaly detection, and pattern recognition. Although DL excels in feature learning, it typically requires extensive training data and has poor interpretability. Conversely, traditional ML methods can operate with smaller datasets and offer more transparent decision-making processes. Therefore, selecting an appropriate ML strategy in practical applications should consider data availability, task complexity, and the need for model interpretability.
@@ VI-A Sleep Staging
The Fpz-Cz channel has been extensively utilized in sleep studies due to its efficiency in detecting wakefulness and REM stages, as well as its ease of use [115, 116, 117]. Wei et al. employed Fpz-Cz channel EEG data from the Sleep-EDFx database, combining contextual scale maps and Co-ScaleNet to achieve high accuracy and recall in automatic sleep staging [115]. Their method proved robust across multiple datasets, including those from healthy individuals, patients with depression and other public datasets, particularly in the N1 stage. The semi-supervised learning model SHNN was proposed for automatic sleep staging using Fpz-Cz channel EEG data from the Sleep-EDFx dataset. This model demonstrated high classification accuracy and consistency, especially in light and deep sleep stages [27]. Studies by [118, 28] used single-channel EEG data from the Fpz-Cz channel for automated sleep staging, exploring sleep staging directly from EEG signals using temporal context information. Eldele et al. proposed an attention-based DL architecture, AttnSleep, to classify sleep stages using single-channel EEG signals, whose architecture relies on a multi-resolution CNN and adaptive feature recalibration [106]. An efficient single-channel EEG analysis model, designed for sleep stage scoring on the Fpz-Cz channel and based on CNN and LSTM, was developed in [104]. An automatic sleep staging system leveraging local extreme value statistical behavior for feature extraction and a residual attention model-based method for accurate EEG sleep staging have been developed, both focusing on the Fpz-Cz channel [119, 120].
The C4-A1 and Cz-A1 channels have demonstrated effective performance for sleep staging tasks on certain datasets. In the SHHS dataset, the C4-A1 channel shows a clear advantage by capturing more information when processing complex EEG signals, thereby improving classification accuracy [118, 121]. The Cz-A1 channel, known for providing stable signals, helps reduce misclassification due to eye movements and its application across multiple datasets has yielded consistent and high-precision classification results. The local extreme value statistical method applied on the C4-A1 channel validated its effectiveness for automatic sleep staging [119], while subsequent experiments with the Cz-A1 channel demonstrated excellent classification performance and advantages in signal stability and eye movement interference reduction [106]. Furthermore, Supratak et al. proposed the TinySleepNet model, which employs DL technology for efficient single-channel EEG analysis. Their study utilized the Cz-A1 channel to verify its efficiency and versatility across multiple datasets. The TinySleepNet model achieves performance comparable to or better than existing methods while reducing the number of parameters and computing resources required [104].